OpenStudy (anonymous):

Derek works 43 hours at a rate of pay of $15 per hour. He receives double pay over 40 hours. What is his gross pay? A. $800 B. $720 C. $780 D. $690

OpenStudy (anonymous):

15 × 40?

OpenStudy (anonymous):

what is 15*40=____

OpenStudy (anonymous):

600

OpenStudy (anonymous):

@bvbarmy17 do 15×40= then 30×3= and add those together

OpenStudy (anonymous):

690 so d